OpenStudy (anonymous):

A box contains 7 red balls. 5 white balls and 3 green balls. In how many ways can we select 3 balls such that- a. They are all red? b. They are all white? c. They are all green? d. They are different colors? e. 2 are red and 1 is white? f. Exactly 2 are white? g. Exactly 1 is green h. None is green?

OpenStudy (sirm3d):

(a) choose 3 red balls among 7. that's combination \(\binom{7}{3}\)

OpenStudy (anonymous):

that is the final answer?

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k\end{matrix}\right)=\frac{n!}{k!(n-k)!}\]